OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#14007 — TH2 <> MAD & GSW <> MAD
Scheduled Maintenance Report for Network & Infrastructure
Completed
Good news:
The link th2-1 n7 <> mad-5-a72 was mounted some minutes ago!

On July 13th 2015 at 2:52:36 p.m. CEST mad-5-a72 Ebra: %LINEPROTO-5-UPDOWN: Line protocol on Interface Ethernet50/1 (I:th2-1-a9:XXX:XXX:PoX), changed state to up

So we will be able to prod TH2 <> MAD and GSW <> MAD, all in 100G of course. :)

We will validate the physical part and mount the ospf!

Update(s):

Date: 2015-07-13 15:26:20 UTC
It is in production on the 100G! We are monitoring and the 10G will be left in backup for some days.

Date: 2015-07-13 15:25:04 UTC
mad-5-a72#sh inter Ethernet50/1-4 transceiver
If device is externally calibrated, only calibrated values are printed.
N/A: not applicable, Tx: transmit, Rx: receive.
mA: milliamperes, dBm: decibels (milliwatts).
Bias Optical Optical
Temp Voltage Current Tx Power Rx Power
Port (Celsius) (Volts) (mA) (dBm) (dBm) Last Update
----- --------- -------- -------- -------- -------- -------------------
Et50/1 36.93 3.34 56.15 0.27 -0.97 0:00:01 ago
Et50/2 36.93 3.34 54.77 -0.20 -0.26 0:00:01 ago
Et50/3 36.93 3.34 51.70 0.36 -0.62 0:00:01 ago
Et50/4 36.93 3.34 62.35 -0.13 -0.62 0:00:01 ago
mad-5-a72#sh inter Ethernet49/1-4 transceiver
If device is externally calibrated, only calibrated values are printed.
N/A: not applicable, Tx: transmit, Rx: receive.
mA: milliamperes, dBm: decibels (milliwatts).
Bias Optical Optical
Temp Voltage Current Tx Power Rx Power
Port (Celsius) (Volts) (mA) (dBm) (dBm) Last Update
----- --------- -------- -------- -------- -------- -------------------
Et49/1 37.88 3.32 56.23 0.30 -2.52 0:00:00 ago
Et49/2 37.88 3.32 54.62 -0.43 -1.65 0:00:00 ago
Et49/3 37.88 3.32 61.50 -0.42 -2.17 0:00:00 ago
Et49/4 37.88 3.32 61.14 -0.27 0.12 0:00:00 ago

Looks good!
Posted Jul 13, 2015 - 15:24 UTC